Q2a. If the probability of passing Accountancy class is 60 percent, Marketing course is 70 percent and probability of passing both Accountaney and Marketing is 35 percent, then what is the probability of passing at least marketing or Accountancy and or both courses? (Show working) 4 points Q2b. Suppose there are 8 tennis balls in a basket each of a different color. Find the numbers of ways in which you can combine and permute the tennis balls if you select 4 tennis balls at a time (Show working). 4 points 21 Page
Answer #1

Q2.a) Let's say probability of passing Accountancy = p(A) = 60%

Probability of passing management = p(M) = 70%

Probability of passing both courses = p(A and M) = 35%

Probability of passing atleast a course = p(A or M) = x

Then by using the probability rule of

p(A or M) = p(A) + p(M) - p(A and M)

p(A or M) = 60% + 70% - 35%

p(A or M) = 95%

Therefore probability of passing either of the course or both is 95%.

Q2.b)

total number of balls = n= 8

No. Of balls to be selected at a time = 4

Therefore total no. Of combinations that can be possible for picking up 4 balls out of 8 balls at a time = 8!/(8-4)!4!

i.e. 40,320/(24×24) = 40,320/576 = 70

Therefore total no. of combinations that can be possible for picking up 4 balls out of 8 different colour balls is 70.